What is the equation of a line with slope -7 that passes through the point (3,2)?

Answer by ankor@dixie-net.com(22740)

What is the equation of a line with slope -7 that passes through the point (3,2)?
:
Use the point/slope form. y - y1 = m(x - x1), where
y1 = 2
x1 = 3
m = -7
y - 2 = -7(x - 3)
distribute -7
y - 2 = -7x + 21
add 2 to both sides
y = -7x + 21 + 2
y = -7x + 23